# LOD-GS: Level-of-Detail-Sensitive 3D Gaussian Splatting for Detail Conserved Anti-Aliasing

### v2

@@ 1 Introduction
Novel view synthesis (NVS) plays a crucial role in the fields of computer vision and computer graphics. Advanced NVS techniques significantly enhance applications in virtual reality, digital modeling, and embodied AI. A notable milestone in NVS is the Neural Radiance Field (NeRF) [27], which represents a 3D scene using a multi-layer perceptron (MLP) and optimizes this MLP through volume ray marching [26, 17, 5] and gradient descent based on multi-view inputs. In comparison to NeRF [27], recent advancements in 3D Gaussian Splatting (3DGS) [14] offer an alternative approach by representing a 3D scene with a collection of 3D Gaussians. This method benefits from an efficiently implemented CUDA rasterizer, which facilitates real-time rendering and efficient training processes. These advantages position 3DGS as a competitive alternative to NeRF.
@@
Figure 1: (a) 3DGS treats a 3D Gaussian primitive (in blue) uniformly across different views and applies a dilation operator (in red) before rendering. (b) Mip-Splatting applies the fixed 3D smoothing filter (in yellow) to the primitive across all views. (c) LOD-GS applies different filters (in yellow and orange) based on the sampling rates of the views. Both Mip-Splatting and our LOD-GS utilize an EWA filter (in red) before rendering. The lack of filtering and the use of a dilation operator in 3DGS result in erosion and dilation effects. The fixed 3D smoothing filter in Mip-Splatting leads to over-smoothing and under-filtering.
@@
However, 3DGS also faces aliasing problems [7] like NeRF. 3DGS is able to achieve good novel view synthesis performance when the training and testing views share roughly the same sampling rate to the observed scene, defined as the ratio of focal length to camera distance. So the aliasing problem of 3DGS is unobvious because existing datasets usually capture a scene from nearly the same distance with the same focal length. For 3DGS models trained on these single-scale datasets, they tend to generate degraded results or aliased renderings when test views zoom in or out significantly. The main reason for this problem is that the 3DGS model is fixed after training. It does not adjust its appearance according to the sampling rate. This limitation leads to the aliasing and degraded results.
@@
In this paper, we are inspired by the Mipmap technique [41], which is used in computer graphics rendering pipelines to tackle the aliasing. Mipmap involves a collection of progressively downsampled textures, enabling the program to choose the appropriate resolution based on the camera’s sampling rate——a technique known as pre-filtering. This method is grounded in the Nyquist-Shannon Sampling Theorem [29, 34] and its concept can be generalized as Level of Detail (LOD) [24, 40, 15, 8], which indicates that the appearance of an object changes with the sampling rate for the scene.
@@
We adopt the Mipmap technique into the 3D Gaussian Splatting framework [14] to represent the pre-filtered radiance field with different levels of detail using multiscale images as input. Because the vanilla 3DGS is not sensitive to the sampling rate, training it with multi-scale images can lead to ambiguity in optimization, making rendering results blur and lack of detail. In this paper, we propose a level-of-detail-sensitive 3DGS framework, named LOD-GS, to effectively sense the change in sampling rate and train the pre-filtered radiance field from pre-filtered images. We propose to add a set of basis functions on each Gaussian primitive to learn the appearance change across different sampling rates. These basis functions take the sampling rate as input and predict the filter size and appearance change for each primitive as illustrated in Figure 1 and Figure 3. The filtered primitive will be splatted into 2D screen space and perform the Elliptical Weighted Average (EWA) filtering [55] to further improve anti-aliasing ability of our method. Besides the framework, we re-render the synthetic dataset used in NeRF [27] from three different camera distances to simulate the changes in sampling rate caused by varying camera distances. The evaluation is carried out on both public datasets [27, 2, 3] and our newly collected dataset. Experiment results demonstrate that our method can disentangle training views with different sampling rates and learn a pre-filtered radiance field from these views, achieving detail-conserved and aliasing-free rendering at the same time.
@@
  • We extend the original NeRF Synthetic Dataset [27] by incorporating rendering views from different camera distances, providing a more comprehensive evaluation of zoom-in and zoom-out effects in neural radiance fields.
@@ 2 Related Work
Novel View Synthesis    Novel View Synthesis (NVS) is a fundamental vision task aimed at generating plausible renderings from new camera positions based on a set of input images and their corresponding camera positions. The introduction of Neural Radiance Fields (NeRF) [27] has significantly changed the approach to solving the NVS task. NeRF formulates Novel View Synthesis as a volume rendering and optimization problem. Nearly all subsequent works adhere to this framework. However, due to the frequent querying of the MLP during training and inference, vanilla NeRF [27] is quite slow, requiring over ten hours for training per scene and dozens of seconds to render a new view. Following works have replaced the full implicit representation of NeRF with feature grid-based representations to accelerate training and inference speeds [6, 10, 23, 35, 28]. Some approaches also explore the point-based neural representations [44, 36, 16, 32, 49, 53]. The emergence of 3D Gaussian Splatting [14], which represents the scene as a collection of 3D Gaussians, has removed the MLP from the radiance field representation. This specially designed MLP-free framework, combined with efficient implementation, allows 3DGS to achieve real-time rendering speeds and perform efficient training. Due to these advantages, 3DGS is beginning to replace NeRF in many applications. Vanilla 3DGS is struggle to tackle the aliasing, this paper focuses on enhancing the anti-aliasing capability of 3DGS to achieve aliasing-free and detail conserved rendering.
@@
Anti-Aliasing of Neural Radiance Field    Neural rendering methods integrate pre-filtering techniques to mitigate aliasing [54, 2, 3, 4, 12, 50]. Mip-NeRF [2] is the first work to address the aliasing problem in NeRF by proposing Integrated Positional Encoding (IPE), which filters out high-frequency components when the camera’s sampling rate is low. Following works [12, 3] speed up the training of Mip-NeRF and extend it to handle unbounded scenes. Due to the differences in rendering methods between NeRF and 3DGS, the anti-aliasing strategy used in NeRF cannot be directly applied to 3DGS. Several methods have recently been proposed to address the aliasing problem in Gaussian Splatting [50, 45, 18, 21]. Multiscale-3DGS (MSGS) [45] represents a 3D scene as a collection of 3D Gaussians of varying levels, selecting only a subset of these levels for rendering based on the target resolution to overcome aliasing. Mip-Splatting [50] introduces the 3D smoothing filter and the 2D mip-filter to remove high-frequency components from the Gaussians. Mipmap-GS [18] utilizes a two stage training and generated pseudo ground truth to adapt to a specific resolution. Analytic-Splatting [21] proposes an approximate method to compute the integration of Gaussian, effectively tackling the aliasing problem. However, all these methods have their limitations. The explicit level representation in MSGS complicates the training and increase the number of primitives. The 3D filter in Mip-Splatting remains fixed after training regardless of the sampling rate of camera, leading to issues of under-filtering and over-smoothing as illustrated in Figure 2. Mipmap-GS needs re-training once the sampling rate changes. For Analytic-Splatting, the introduction of integration slows down its training and rendering. In this paper, we propose a level-of-detail-sensitive filtering strategy to effectively achieve aliasing-free rendering.
Additional Parameter Control for Gaussian Primitive    We aim to adjust the degree of filtering applied on Gaussians according to the value of sampling rate. Therefore, we review how existing methods modify the parameters of 3DGS using an additional input variable in this section. Aside from storing the parameter changes at all time steps [25], the remaining approaches can be broadly classified into three categories. The methods in the first category use MLPs to model parameter changes [42, 47, 13], which has been explored in NeRF area [11, 20, 30, 31, 33, 37, 43, 51]. For instance, 4DGaussians [42] queries a learnable spatial-temporal feature grid using Gaussian positions, then inputs this feature into a neural network to predict spatial shifts and deformations of the primitive. Another category represents a scene as a collection of high-dimensional Gaussians. The additional input is treated as a new dimension of the Gaussian. During rendering, given the value of the input variable on this new axis, the conditioned 3D Gaussians are computed before rendering [48, 9]. The main drawback of this design is that it can easily generate large amounts of high-dimensional Gaussians, leading to overfitting and increased model size. In addition to the above two categories, a different approach to modeling parameter changes is to add a set of learnable basis functions to each Gaussian primitive to capture variations over the input variable [22, 46]. In this paper, we model the filter change of 3DGS using a set of basis functions that take the camera’s sampling rate as input. This design helps us predict the most appropriate filter for aliasing-free rendering in an efficient manner.
@@ 3 Preliminaries
Kerbl et al. [14] utilize learnable 3D Gaussian primitives to represent 3D scenes and render different views using a differentiable volume splatting rasterizer. In 3DGS, each 3D Gaussian primitive is parameterized using a 3D covariance matrix $\boldsymbol{\Sigma}$ and a distribution center $\mathbf{p}_{k}$:
@@
During optimization, the covariance matrix $\boldsymbol{\Sigma}$ is factorized into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as $\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$ to ensure its positive semidefiniteness. To obtain the rendering results of 3D Gaussians from a specific view, the 3D Gaussian is first projected to a 2D splat in screen space using the view matrix $\mathbf{W}$ and an affine approximated projection matrix $\mathbf{J}$ as illustrated in [55]:
@@ 3.2 Nyquist-Shannon Sampling Theorem
The rendering process of 3D Gaussian Splatting involves sampling. It uses discrete pixels to sample 3D Gaussians, ultimately producing a 2D rendering result from this 3D representation. Consequently, the rendering of 3D Gaussian Splatting also faces the challenge of aliasing. Corresponding to the two approaches mentioned earlier, the first method to address aliasing in 3DGS rendering is Super Sampling [38]. This technique renders images at a higher resolution—effectively increasing the sampling rate—and then pools the high-resolution results to generate anti-aliased low-resolution images. However, the high resolution used in Super Sampling leads to additional computational burden, resulting in slower rendering speeds. The second approach applies filtering to 3D Gaussians before rendering, ensuring their frequency remains below the Nyquist limit. We adopt this method for efficient rendering. However, excessive filtering can cause signal loss, leading to overly smooth results that lack fine details. In Section 4, we will explain in detail how our method selects the appropriate filter for each Gaussian primitive.
@@
NeRF [27] 29.90 32.13 33.40 29.47 31.23 0.938 0.959 0.973 0.962 0.958 0.074 0.040 0.024 0.039 0.044
MipNeRF [2] 32.63 34.34 35.47 35.60 34.51 0.958 0.970 0.979 0.983 0.973 0.047 0.026 0.017 0.012 0.026
Tri-MipRF [12] 32.65 34.24 35.02 35.53 34.36 0.958 0.971 0.980 0.987 0.974 0.047 0.027 0.018 0.012 0.026
3DGS [14] 28.79 30.66 31.64 27.98 29.77 0.943 0.962 0.972 0.960 0.960 0.065 0.038 0.025 0.031 0.040
Mipmap-GS [19] 28.79 30.67 31.66 28.00 29.78 0.943 0.962 0.973 0.961 0.960 0.065 0.038 0.025 0.031 0.040
Multiscale-3DGS[45] 33.36 27.15 21.41 17.61 24.88 0.969 0.951 0.875 0.764 0.890 0.031 0.032 0.067 0.126 0.064
Mip-Splatting [50] 32.81 34.49 35.45 35.50 34.56 0.967 0.977 0.983 0.988 0.979 0.035 0.019 0.013 0.010 0.019
Analytic-Splatting [21] 33.22 34.92 35.98 36.00 35.03 0.967 0.977 0.984 0.989 0.979 0.033 0.019 0.012 0.010 0.018
LOD-GS (ours) 32.90 34.88 36.43 37.27 35.37 0.966 0.977 0.984 0.990 0.979 0.036 0.019 0.012 0.008 0.019
w/o LOD filter 31.88 34.09 35.86 35.73 34.39 0.961 0.975 0.984 0.989 0.977 0.042 0.021 0.013 0.010 0.022
w/o EWA filter 32.63 34.39 35.66 35.76 34.61 0.966 0.977 0.983 0.987 0.978 0.038 0.020 0.013 0.010
Table 1: Multi-scale Training and Multi-scale Testing on the Blender dataset [27]. Our approach demonstrates state-of-the-art performance across most metrics and is highly competitive with existing GS-based methods specifically designed for anti-aliasing, such as Multiscale-3DGS [45], Mip-Splatting [50], Mipmap-GS [18], and Analytic-Splatting [21].
@@ 4.1 Level-of-Detail-Sensitive 3D Filter
Applying zoom-in and zoom-out operations on scenes reconstructed by 3DGS can lead to noticeable visual artifacts, including erosion and dilation [50]. To solve these artifacts, Mip-Splatting proposes 3D Smoothing Filter as follows :
@@
In the Level of Detail (LOD) concept [24, 40, 15, 8], an object’s texture resolution adapts based on the sampling rate. Inspired by this idea, we design a learnable framework that takes the sampling rates of Gaussian primitives as input and outputs appropriate 3D filters. In other words, the filter size of each primitive dynamically adjusts according to the sampling rate. To minimize the increase in computational workload of training and inference, we design a learnable Gaussian Mixture Model (GMM) module instead of using a MLP for each primitive to predict the suitable filter size and appearance change. The added GMM module can be expressed using the following equation:
@@
$\mathcal{G}_{k}(\mathrm{x})=(\alpha_{k}+\mathcal{F}_{\alpha}(\nu))e^{-\frac{1}{2}(\mathrm{x}-\mathbf{p}_{k})^{T}(\Sigma_{k}+\mathcal{F}_{s}(\nu))(\mathrm{x}-\mathbf{p}_{k})}$ (7)
where $\nu$ is the sampling rate of the Gaussian primitive. The sampling rate from one camera to all the primitives can be computed efficiently in $\mathcal{O}(K)$. This allows us to perform this computation in each iteration, rather than updating the sampling rate every 100 iterations, as done in Mip-Splatting [50]. This design makes each Gaussian primitive sensitive to changes in the sampling rate. As a result, the scale and opacity of the primitive can be adjusted accordingly. The following experimental results demonstrate that our proposed method can better bake images from different sampling rates into a single radiance field, generating aliasing-free rendering results while preserving details.

### v3

@@ 1 Introduction
Novel view synthesis (NVS) plays a crucial role in the fields of computer vision and computer graphics. Advanced NVS techniques significantly enhance applications in virtual reality, digital modeling, and embodied AI. A notable milestone in NVS is the Neural Radiance Field (NeRF) [mildenhall2020nerf], which represents a 3D scene using a multi-layer perceptron (MLP) and optimizes this MLP through volume ray marching [max1995optical, levoy1990efficient, brebin1998volume] and gradient descent based on multi-view inputs. In comparison to NeRF [mildenhall2020nerf], recent advancements in 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] offer an alternative approach by representing a 3D scene with a collection of 3D Gaussians. This method benefits from an efficiently implemented CUDA rasterizer, which facilitates real-time rendering and efficient training processes. These advantages position 3DGS as a competitive alternative to NeRF.
@@
Figure 1: (a) 3DGS treats a 3D Gaussian primitive (in blue) uniformly across different views and applies a dilation operator (in red) before rendering. (b) Mip-Splatting applies the fixed 3D smoothing filter (in yellow) to the primitive across all views. (c) LOD-GS applies different filters (in yellow and orange) based on the sampling rates of the views. Mip-Splatting utilizes 2D Mip filter (in red) and LOD-GS uses EWA filter (in red) before rendering. The lack of filtering and the use of a dilation operator in 3DGS result in erosion and dilation effects. The fixed 3D smoothing filter in Mip-Splatting leads to over-smoothing and under-filtering.
@@
However, 3DGS also faces aliasing problems [crow1977aliasing] like NeRF. 3DGS is able to achieve good novel view synthesis performance when the training and testing views share roughly the same sampling rate to the observed scene, defined as the ratio of focal length to camera distance. So the aliasing problem of 3DGS is unobvious because existing datasets usually capture a scene from nearly the same distance with the same focal length. For 3DGS models trained on these single-scale datasets, they tend to generate degraded results or aliased renderings when test views zoom in or out significantly. The main reason for this problem is that the 3DGS model is fixed after training. It does not adjust its appearance according to the sampling rate. This limitation leads to the aliasing and degraded results.
@@
In this paper, we are inspired by the Mipmap technique [mipmap], which is used in computer graphics rendering pipelines to tackle the aliasing. Mipmap involves a collection of progressively downsampled textures, enabling the program to choose the appropriate resolution based on the camera’s sampling rate——a technique known as pre-filtering. This method is grounded in the Nyquist-Shannon Sampling Theorem [nyquist2009certain, shannon1949communication] and its concept can be generalized as Level of Detail (LOD) [lod-of-3d, weiler2000level, lamar1999multiresolution, de1995hierarchical], which indicates that the appearance of an object changes with the sampling rate for the scene.
@@
We adopt the Mipmap technique into the 3D Gaussian Splatting framework [kerbl3Dgaussians] to represent the pre-filtered radiance field with different levels of detail using multiscale images as input. Because the vanilla 3DGS is not sensitive to the sampling rate, training it with multi-scale images can lead to ambiguity in optimization, making rendering results blur and lack of detail. In this paper, we propose a level-of-detail-sensitive 3DGS framework, named LOD-GS, to effectively sense the change in sampling rate and train the pre-filtered radiance field from pre-filtered images. We propose to add a set of basis functions on each Gaussian primitive to learn the appearance change across different sampling rates. These basis functions take the sampling rate as input and predict the filter size and appearance change for each primitive as illustrated in Figure 1 and Figure 3. The filtered primitive will be splatted into 2D screen space and perform the Elliptical Weighted Average (EWA) filtering [zwicker2001ewa] to further improve anti-aliasing ability of our method. Besides the framework, we re-render the synthetic dataset used in NeRF [mildenhall2020nerf] from three different camera distances to simulate the changes in sampling rate caused by varying camera distances. The evaluation is carried out on both public datasets [mildenhall2020nerf, mip-nerf, mipnerf360] and our newly collected dataset. Experiment results demonstrate that our method can disentangle training views with different sampling rates and learn a pre-filtered radiance field from these views, achieving detail-conserved and aliasing-free rendering at the same time.
@@
  • We extend the original NeRF Synthetic Dataset [mildenhall2020nerf] by incorporating rendering views from different camera distances, providing a more comprehensive evaluation of zoom-in and zoom-out effects in neural radiance fields.
@@ 2 Related Work
Novel View Synthesis    Novel View Synthesis (NVS) is a fundamental vision task aimed at generating plausible renderings from new camera positions based on a set of input images and their corresponding camera positions. The introduction of Neural Radiance Fields (NeRF) [mildenhall2020nerf] has significantly changed the approach to solving the NVS task. NeRF formulates Novel View Synthesis as a volume rendering and optimization problem. Nearly all subsequent works adhere to this framework. However, due to the frequent querying of the MLP during training and inference, vanilla NeRF [mildenhall2020nerf] is quite slow, requiring over ten hours for training per scene and dozens of seconds to render a new view. Following works have replaced the full implicit representation of NeRF with feature grid-based representations to accelerate training and inference speeds [chen2022tensorf, fridovich2022plenoxels, liu2020neural, sun2022direct, muller2022instant]. Some approaches also explore the point-based neural representations [xu2022point, sun2024pointnerf++, lassner2021pulsar, prokudin2023dynamic, yifan2019differentiable, zheng2023pointavatar]. The emergence of 3D Gaussian Splatting [kerbl3Dgaussians], which represents the scene as a collection of 3D Gaussians, has removed the MLP from the radiance field representation. This specially designed MLP-free framework, combined with efficient implementation, allows 3DGS to achieve real-time rendering speeds and perform efficient training. Due to these advantages, 3DGS is beginning to replace NeRF in many applications. However, 3DGS is struggling to tackle the aliasing, this paper focuses on enhancing the anti-aliasing capability of 3DGS to achieve aliasing-free and detail conserved rendering.
@@
Anti-Aliasing of Neural Radiance Field    Neural rendering methods integrate pre-filtering techniques to mitigate aliasing [zhuang2023lod, mip-nerf, mipnerf360, barron2023zip, trimip-nerf, mip-splatting]. Mip-NeRF [mip-nerf] is the first work to address the aliasing problem in NeRF by proposing Integrated Positional Encoding (IPE), which filters out high-frequency components when the camera’s sampling rate is low. Following works [trimip-nerf, mipnerf360] speed up the training of Mip-NeRF and extend it to handle unbounded scenes. Due to the differences in rendering methods between NeRF and 3DGS, the anti-aliasing strategy used in NeRF cannot be directly applied to 3DGS. Several methods have recently been proposed to address the aliasing problem in Gaussian Splatting [mip-splatting, multiscale-3dgs, mipmapgs, analytic-splatting, sa-gs, octreegs]. Mip-Splatting [mip-splatting] introduces the 3D smoothing filter and the 2D mip-filter to remove high-frequency components from the Gaussians. Mipmap-GS [mipmapgs] utilizes a two stage training and generated pseudo ground truth to adapt to a specific resolution. Analytic-Splatting [analytic-splatting] proposes an approximate method to compute the integration of Gaussian, effectively tackling the aliasing problem. However, all these methods have their limitations. Mipmap-GS needs re-training once the sampling rate changes. The introduction of integration in Analytic-Splatting slows down its training and rendering. The 3D filter in Mip-Splatting remains fixed after training, regardless of the camera’s sampling rate. Additionally, the 2D Mip filter cannot sense the existence of the 3D filter, potentially resulting in inappropriate 2D filtering degree. All these factors contribute to the problem of under-filtering and over-smoothing, as illustrated in Figure 2. To solve these problems, we propose a filtering strategy which dynamically adjusts the filtering degree and learns this adjustment from data in an end-to-end manner.
@@ 3 Preliminaries
Kerbl et al. [kerbl3Dgaussians] utilize learnable 3D Gaussian primitives to represent 3D scenes and render different views using a differentiable volume splatting rasterizer. In 3DGS, each 3D Gaussian primitive is parameterized using a 3D covariance matrix $\boldsymbol{\Sigma}$ and a distribution center $\mathbf{p}_{k}$:
@@
During optimization, the covariance matrix $\boldsymbol{\Sigma}$ is factorized into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as $\boldsymbol{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\top}\mathbf{R}^{\top}$ to ensure its positive semidefiniteness. To obtain the rendering results of 3D Gaussians from a specific view, the 3D Gaussian is first projected to a 2D splat in screen space using the view matrix $\mathbf{W}$ and an affine approximated projection matrix $\mathbf{J}$ as illustrated in [zwicker2001ewa]:
@@ 3.2 Nyquist-Shannon Sampling Theorem
The rendering process of 3D Gaussian Splatting involves sampling. It uses discrete pixels to sample 3D Gaussians, ultimately producing a 2D rendering result from this 3D representation. Consequently, the rendering of 3D Gaussian Splatting also faces the challenge of aliasing. Corresponding to the two approaches mentioned earlier, the first method to address aliasing in 3DGS rendering is Super Sampling [wang2022nerfsr-mm]. This technique renders images at a higher resolution—effectively increasing the sampling rate—and then pools the high-resolution results to generate anti-aliased low-resolution images. However, the high resolution used in Super Sampling leads to additional computational burden, resulting in slower rendering speeds. The second approach applies filtering to 3D Gaussians before rendering, ensuring their frequency remains below the Nyquist limit. We adopt this method for efficient rendering. However, excessive filtering can cause signal loss, leading to overly smooth results that lack fine details. In Section 4, we will explain in detail how our method selects the appropriate filter for each Gaussian primitive.
@@
NeRF [mildenhall2020nerf] 29.90 32.13 33.40 29.47 31.23 0.938 0.959 0.973 0.962 0.958 0.074 0.040 0.024 0.039 0.044
MipNeRF [mip-nerf] 32.63 34.34 35.47 35.60 34.51 0.958 0.970 0.979 0.983 0.973 0.047 0.026 0.017 0.012 0.026
Tri-MipRF [trimip-nerf] 32.65 34.24 35.02 35.53 34.36 0.958 0.971 0.980 0.987 0.974 0.047 0.027 0.018 0.012 0.026
3DGS [kerbl3Dgaussians] 28.79 30.66 31.64 27.98 29.77 0.943 0.962 0.972 0.960 0.960 0.065 0.038 0.025 0.031 0.040
Mipmap-GS [mipmapgs] 28.79 30.67 31.66 28.00 29.78 0.943 0.962 0.973 0.961 0.960 0.065 0.038 0.025 0.031 0.040
SA-GS [sa-gs] 30.80 32.67 35.06 35.77 33.58 0.956 0.969 0.980 0.985 0.973 0.056 0.032 0.020 0.014 0.031
Multiscale-3DGS[multiscale-3dgs] 33.36 27.15 21.41 17.61 24.88 0.969 0.951 0.875 0.764 0.890 0.031 0.032 0.067 0.126 0.064
Mip-Splatting [mip-splatting] 32.81 34.49 35.45 35.50 34.56 0.967 0.977 0.983 0.988 0.979 0.035 0.019 0.013 0.010 0.019
Analytic-Splatting [analytic-splatting] 33.22 34.92 35.98 36.00 35.03 0.967 0.977 0.984 0.989 0.979 0.033 0.019 0.012 0.010 0.018
LOD-GS (ours) 32.90 34.88 36.43 37.27 35.37 0.966 0.977 0.985 0.990 0.980 0.035 0.019 0.012 0.008
Table 1: Multi-scale Training and Multi-scale Testing on the Blender dataset [mildenhall2020nerf]. Our approach demonstrates state-of-the-art performance across most metrics and is highly competitive with existing GS-based methods specifically designed for anti-aliasing, such as Mipmap-GS [mipmapgs], SA-GS [sa-gs], Multiscale-3DGS [multiscale-3dgs], Mip-Splatting [mip-splatting], and Analytic-Splatting [analytic-splatting].
@@ 4.1 Level-of-Detail-Sensitive 3D Filter
Applying zoom-in and zoom-out operations on scenes reconstructed by 3DGS can lead to noticeable visual artifacts, including erosion and dilation [mip-splatting]. To solve these artifacts, Mip-Splatting proposes 3D Smoothing Filter as follows :
@@
In the Level of Detail (LOD) concept [lod-of-3d, weiler2000level, lamar1999multiresolution, de1995hierarchical], an object’s texture resolution adapts based on the sampling rate. Inspired by this idea, we design a learnable framework that takes the sampling rates of Gaussian primitives as input and outputs appropriate 3D filters. In other words, the filter size of each primitive dynamically adjusts according to the sampling rate. To minimize the increase in computational workload of training and inference, we design a learnable Gaussian Mixture Model (GMM) module instead of using a MLP for each primitive to predict the suitable filter size. The added GMM module can be expressed using the following equation:
@@
$\mathcal{G}_{k}(\mathrm{x})=e^{-\frac{1}{2}(\mathrm{x}-\mathbf{p}_{k})^{T}(\Sigma_{k}+\mathcal{F}_{s}(\nu))(\mathrm{x}-\mathbf{p}_{k})}$ (7)
where $\nu$ is the sampling rate of the Gaussian primitive. The sampling rate from one camera to all the primitives can be computed efficiently in $\mathcal{O}(K)$. This allows us to perform this computation in each iteration, rather than updating the sampling rate every 100 iterations, as done in Mip-Splatting [mip-splatting]. This design makes each Gaussian primitive sensitive to changes in the sampling rate. As a result, the scale and opacity of the primitive can be adjusted accordingly. The following experimental results demonstrate that our proposed method can better bake images from different sampling rates into a single radiance field, generating aliasing-free rendering results while preserving details.